Chemist mixes 200 ml of solution that is 15% salt with 300 ml that is 40% salt. What percent salt is resulting solution?
Answer by nerdybill(7384)

Chemist mixes 200 ml of solution that is 15% salt with 300 ml that is 40% salt. What percent salt is resulting solution?
Let x = percent salt of resulting solution
then
.15(200) + .40(300) = x(200+300)
30 + 120 = x(500)
150 = x(500)
150/500 = x
0.30 = x
or
30% = x